Operation Summary for October 20, 2010 - October 22, 2010 |
Wednesday Oct. 20 The day shift began with Operations monitoring the Tevatron (TeV) store 8188 (139E30) and the Recycler stash (135E10), with the Antiproton Source (Pbar) stacking (10E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with NuMI off for target work, with the MiniBooNE and Meson FTBF CALICE T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations halted beam to Pbar at 9 AM for a vault access to install equipment for the T944 JASMIN experiment. Pbar personnel inspected the Pbar target and found damage that required a target replacement. RF technicians accessed Booster to replace two power amplifiers and MiniBooNE personnel accessed MI-12B to fill the LCW system. Operations resumed sending beam to MiniBooNE at 12:35 PM and to T-978 shortly after. During the evening shift, at 5:09 PM, store 8188 aborted due to an upstream quench indication on a TeV HOPS power supply (T:QDD1); EE Support personnel investigated and replaced its quench detection cards. Operators prepared CDF and D0 for collision hall accesses. At 5:38 PM, the Recycler stash was lost; the current theory points to a clearing electrode at MI-30 that was off and allowed trapped ions to induce a half-integer resonance. Operations disabled beam to T-978 at 6:20 PM. Operations reported that QDD1 failed again on another downstream quench indication. EE Support recalibrated the quench protection cards. Pbar personnel continued to work on the target replacement. Thursday Oct. 21 The midnight shift began with EE Support personnel working on T:QDD1, with Pbar personnel replacing their target, with the Recycler Pelletron work continuing, with Operations monitoring the stash (16.2E10) and the Pbar stack (6.5E10), and with MiniBooNE taking beam. During the midnight shift, Operations reported that the target work concluded and Pbar started stacking at 12:41 AM. Operators topped off TeV sector E48 kicker with insulating gas. Operations reported that the MiniBooNE beam permit went down when their LCW system tripped off with no fault indications. Operators consulted with the MiniBooNE personnel and resumed sending beam to the experiment. Operations reported that an unknown Booster RF problem had caused about 15 minutes of downtime. This was followed by MI RF trips, APG-B overdrive faults, which required a reboot of the MI low-level RF system and the dampers. EE Support reported that they had completed their TeV QDD1 work at 3:14 AM. Operations put the TeV through a dry squeeze, but the TeV ramp locked at 90 GeV when coming out of the dry squeeze due to cryo system problems at TeV sector A2 that took down the cryo system beam permit. Operators resolved these problems, put the TeV through another dry squeeze, followed by wet squeezes, and then conducted orbit checks. Operations established beam to T-978 at 6:51 AM. During the day shift, Operations put the TeV into shot setup at 8:57 AM and established store 8194 at 10:01 AM with an average initial luminosity (AIL) of 122.3E30. Unfortunately, the store suffered a CDF SVX abort at 3:06 PM due to some work around the BLM high voltage chassis at CDF that caused a false beam loss indication. The Pelletron came online at 3:21 PM. During the evening shift, Operations reported that the TeV quenched during the orbit check due to an apparent T:QDD1 regulation problem. During recovery, a beam valve in TeV sector A1 was found closed and would not open. Operators prepared the Transfer hall for access to allow Mechanical personnel to begin repairs. Operations disabled the T-978 beam at 6 PM. Friday Oct. 22 The midnight shift began with the TeV Transfer Hall in access, with Operations monitoring the stash (351.1E10) and the Pbar stack (23.9E10), and with MiniBooNE taking beam. During the midnight shift, Mechanical personnel came out of the Transfer Hall at 1 AM. Operations conducted TeV checkout and then put the TeV through dry and wet squeezes. Operations put the TeV into shot setup at 4:25 AM and established store 8198 at 5:32 AM with an AIL of 305.8E30. Operations reported that while trying to bring up T-978, a Meson power supply (M01D) tripped off on a magnet LCW flow switch indication. Operators called in Waterworks personnel to deal with the problem. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Linac personnel to conduct two small jobs today. |
|
More Information Comments and Suggestions |
| last modified 10/22/2010 email Fermilab |
| Security, Privacy, Legal | |